f(x)=x^3-5x^2-14x


f(x)=x^3+4x^2+9x+36


f(x)=x^4+x^3-2x^2+4x-24

please helpp!

Answer by checkley71(8403)

F(X)=X^3-5X^2-14X
F(X)=X(X^2-5X-14)
F(X)=X(X-7)(X+2)
----------------------------------
F(X)=X^3+4X^2+9X+36
F(X)=X^2(X+4)+9(X+4)
F(X)=(X^2+9)(X+4)^2
----------------------------------
F(X)=X^4+X^3-2X^2+4X-24
F(X)=X^3(X+1)-2(X^2-2X+12)